# The Commensal Real-time ASKAP Fast Transient incoherent-sum survey

### v1

@@ Abstract
With wide-field phased array feed technology, the Australian Square Kilometre Array Pathfinder (ASKAP) is ideally suited to search for seemingly rare radio transient sources. The Commensal Real-time ASKAP Fast Transient (CRAFT) Survey Science Project has developed instrumentation to continuously search for fast radio transients (duration $\lesssim$ 1 second) with ASKAP, with a particular focus on finding and localising Fast Radio Bursts (FRBs). Of particular interest are Fast Radio Bursts (FRBs). Since 2018, the CRAFT survey has been searching for FRBs and other fast transients by incoherently adding the intensities received by individual ASKAP antennas, and then correcting for the impact of frequency dispersion on these short-duration signals in the resultant incoherent sum (ICS) in real-time. This low-latency detection enables the triggering of voltage buffers, which facilitates the localisation of the transient source and the study spectro-polarimetric properties at high time resolution. Here we report the sample of $43$ FRBs discovered in this CRAFT/ICS survey to date. This includes $22$ FRBs that had not previously been reported: $16$ FRBs localised by ASKAP to $\lesssim 1$ arcsec and $6$ FRBs localised to $\sim 10$ arcmin. Of the new arcsecond-localised FRBs, we have identified and characterised host galaxies (and measured redshifts) for $11$. The median of all $30$ measured host redshifts from the survey to date is $z=0.23$. We summarise results from the searches, in particular those contributing to our understanding of the burst progenitors and emission mechanisms, and on the use of bursts as probes of intervening media. We conclude by foreshadowing future FRB surveys with ASKAP using a coherent detection system that is currently being commissioned. This will increase the burst detection rate by a factor of approximately ten and also distance to which ASKAP can localise FRBs.
@@ 1 INTRODUCTION
Here we summarise the results of this CRAFT/ICS survey and describe the bursts detected. The host galaxies for many of the FRBs discovered are displayed in Figures 1, 2, and 3. In Section 2, we describe the instrumentation and methods used to undertake the searches. We motivate and describe the observing strategies we undertook to find the FRB sample in Section 3 and our multi-wavelength follow up in Section 4. In Section 5, we present the fast radio bursts sample, being 43 FRBs, of which 37 have measured positions with precisions $\lesssim 1^{\prime\prime}$, of which 30 host galaxies have been identified. We assess survey performance in Section 6, and in Section 7 we provide a review of the main scientific findings and outcomes of the searches. In Section 8, we briefly describe the plan for new FRB search systems for ASKAP.
@@
Figure 1: Montage of FRB host galaxies for ASKAP-localised FRBs with firm associations. We present the FRBs in increasing redshift. In the top left corner, we show the FRB name and redshift. In the top right corner, we list the telescope and observing band of the image. The angular and physical scale at the host redshift is shown in the lower left corner. The 1-$\sigma$ localisation region of the FRB is given by the black ellipse.
@@ 2.1 Hardware
Each PAF receiver comprises $188$ dipole elements arranged in a chequerboard pattern. The signal from each dipole element is transmitted from the PAF to the MRO control building via radio frequency over fibre systems Beresford et al. (2017). Observing is undertaken in one of three bands (bands 1,2, and 3, with 1 being the lowest and 3 the highest). Within each band, it is possible to tune the observing frequencies best suited to the science goals. The signal from each element is then digitised and channelised in a digital receiver Brown et al. (2014). The signals are channelised using an oversampled polyphase filter into coarse $1$ MHz channels by a factor of 32/27 Tuthill et al. (2012).
@@ 2.2 FRB Searches
Despite the MRO having low levels of nearby interference, RFI is present in the searches. The largest sources of interference are satellite interference from global navigation systems in band-2 observations and ionospheric ducting of mobile phone signals (base towers and handsets) in band 1. Data were also affected by interference associated with lightning and electrical storms (even when storms were more than $100$ km from the observatory). Real time radio frequency interference (RFI) flagging is conducted using a number of heuristics. Frequency channels with abnormally high kurtosis are set to zero, while remaining channels are normalised to a mean of zero and unit standard deviation. Impulsive RFI is removed by subtracting the frequency integrated signal at zero dispersion measure (DM). This is applied dynamically - additionally, a hard-coded list of known RFI channels is maintained, and these frequencies are zeroed for all observations.
@@
The searches produce a stream of raw candidates (time, signal-to-noise ratio, dispersion measure, width) which are broadcast over User Datagram Protocol. A burst (either RFI or a bona-fide FRB) produces candidates over a region of candidate parameter space. A density-based spatial clustering algorithm Ester et al. (1996) is used to group raw candidates. This is important as it allows RFI (which typically has the greatest S/N at low DM) to be distinguished from FRBs. To ensure that the real-time clustering algorithm was working properly, raw candidates were also inspected off-line after being clustered using a friends-of-friends algorithm Huchra & Geller (1982). The search latency ranges from a few hundred milliseconds to a few seconds. The time resolution of searches varied over the course of the survey, between $860\,\mu$s and $1.7$ ms, as system performance was optimised. There is a trade-off between the number of antennas that can be ingested and the time resolution. For the majority of the searching, a time resolution of $1.182$ ms was chosen, which allowed the incoherent sum of intensities from $\sim$24 antennas to be formed.
@@ 4.2 Host Associations
We revisit this assumption and suggest an updated formulation for this offset prior. Figure 6 shows the normalised PDF of the angular offsets $\theta$ relative to $\phi$ for 17 FRBs. This FRB sample was restricted to the subset of FRBs presented in Bhandari et al. (2022) with $P_{\rm host}>0.85$ (using their PATH analysis) and also the $11$ new FRBs presented in this manuscript with a secure association (see section 5.2). We also include the ASKAP-localised CHIME repeating source FRB 20201124A (Fong et al., 2021). Approximately half of the sample shows $\theta/\phi<1$ with the remainder exhibiting a tail to $\theta/\phi\approx 5$.
Overplotted on the observed distribution are the exponential priors for host galaxy offset with scale lengths of $\phi$ and $\phi/2$. These have been convolved with the reported uncertainties in the FRB localisations. They have also been weighted by a geometrical factor ($2\pi\theta$) which disfavours low offsets [^4] We note that Figure 11 of Aggarwal et al. (2021) failed to include this factor.. Clearly, the exponential with scale-length of $\phi$ is disfavoured while the data are reasonably well-described by the smaller scale length ($\phi/2$). We advocate adopting this new prior for future work on FRBs with PATH, and we utilise it throughout the manuscript.
Reanalysing all of the FRBs presented in Bhandari et al. (2020b), we find few changes in the posterior probabilities. Most of them were previously $P_{\rm host}>0.95$ and the values increased towards 1. The only notable changes were significant increases in $P_{\rm host}$ for FRB 20181112A and FRB 20191001A. The former FRB is associated with a $z=0.4755$ which has a $z=0.3674$ galaxy in the foreground Prochaska et al. (2019) The latter FRB is associated with a $z=0.2340$ galaxy which is separated by $\approx 5$” from a galaxy at a similar redshift ($z=0.2339$ Bhandari et al., 2020a). In Tables 3 to 6 in the Appendix we present PATH probabilities for nearby host galaxies for all ASKAP-localised FRBs.
Figure 6: Comparison of PATH prior distributions to FRB offsets obtained through CRAFT/ICS observations for a subset of CRAFT FRBs.
@@ 4.3 Optical spectroscopy
We obtained spectroscopic observations of the probable host galaxy candidates that lacked archival data, both to determine the redshift but also to study the global properties of the hosts. These were primarily obtained with FORS2 or X-shooter Vernet et al. (2011) on the VLT, or LRIS or the Deep Imaging Multi-Object Spectrograph (DEIMOS; Faber et al. 2003) on Keck. Brighter host galaxies were observed with the Goodman Spectrograph on the Southern Astrophysical Research Telescope (SOAR, Clemens et al., 2004). Spectroscopic data reduction is described in detail in Glowacki et al. (2023), Gordon et al. (2023), Ryder et al. (2023), and A. Muller et al. (in prep).
@@ 5 FAST RADIO BURST DISCOVERIES
Table 1 in the Appendix summarises the key properties of the FRBs. Table 2 in the Appendix lists the positions of the FRBs, derived interferometrically (where possible) and with the multi-beam localisation method Bannister et al. (2017). Dedispersed dynamic spectra of the bursts from the ICS search data stream, arranged from lowest to highest dispersion measure, can be found in Figures 19 to 23 in the appendix. Even with the relatively low time and frequency resolution of the search data stream the bursts show a variety of spectral and temporal morphologies. Their high-time-resolution properties are reported elsewhere (D. Scott et al., in prep.). We explore the spectral modulation of the bursts in Section 6.9.
@@ 5.2 New localised FRBs
We summarise the properties and localisations of a few other FRBs discovered in the ICS searches, though detailed analyses of the bursts and their host galaxies is deferred to subsequent studies.
@@ 5.2.14 FRB 20240304A
The FRB was detected commensally with EMU. VLT/FORS2 imaging has been taken and a host galaxy identified (See Figure 3). Spectroscopic observations have not yet been attempted.
@@ 6 SURVEY PERFORMANCE
Given the commensal nature of the survey, we assess detection rates as a function of changing observation parameters. Through to the end of 2023 (MJD 60309), ASKAP has observed in ICS mode for a total of 15 324.715324.715\,324.715 324.7 hr, for an average on-sky efficiency of $32\%$. As a commensal survey, this has resulted in a variety of pointing directions and frequency configurations. Full logging information was available only for 12 701.712701.712\,701.712 701.7 hr of this, due to an early logging error. This corresponds to the latter 30 FRBs detected until the end of 2023. Furthermore, the current version (‘v3’; 28 FRBs) of the FREDDA FRB detection algorithm has been used only from April 2020 — the initial version (‘v1’; 7 FRBs) had slightly reduced sensitivity to high-DM FRBs, while ‘v2’ (2 FRBs) had a bug that reported incorrect S/N values (Hoffmann et al., 2024). Therefore, some analyses below will be applied to a subset of the data. This will be noted in each case.
@@ 6.2 Modelling frequency dependence
We model the frequency-dependence of the detection rate $R(\nu)$ to be
$R(\nu)$ $=$ $T(\nu)F(\nu)W(\nu)N_{\rm ant}^{0.75}\nu^{\alpha}\,,$ (1)
@@ 6.4 Source counts analysis
The observed source counts distribution is also a good diagnostic tool for biases in the search pipeline: human vetting and/or RFI mitigation algorithms may reject candidates that are close to the nominal detection threshold, or well above it (Macquart & Ekers, 2018b). We have performed Kolmogorov-Smirnov tests (Kolmogorov, 1933; Smirnov, 1948) for consistency with a pure $\alpha=-1.50$ power-law on ICS FRBs, with p-values quoted in Figure 11, and find no strong statistical evidence for a deviation. The same conclusion is reached when dividing the sample into FRBs detected in low, medium, and high frequency ranges.
@@ 6.9 Burst modulation
Many of the ASKAP FRBs show spectral modulation: intensity variation in frequency that could either be diffractive scintillation or intrinsic to the emission mechanism. This was identified in the fly’s eye surveyShannon et al. (2018). We investigated if there is any correlation between modulation and dispersion measure using the FRBs in the search data stream. We choose to investigate modulation and dispersion measure to consider both the fly’s eye and ICS FRBs. We calculate the spectral modulation index to be
@@
where $\sigma_{\rm FRB}^{2}$ is the variance on pulse and $\sigma_{n}^{2}$ is the variance of (thermal) noise measured in a segment of data of equal temporal duration close to the FRB. While FRBs with DM $\lesssim 500\,$pc cm-3show a variety of spectral modulation, those with DM $>800\,$pc cm-3show little spectral modulation. It is possible that spectral modulation is being quenched at higher DM. Studies of a larger sample of bursts in higher spectral and time resolution (such as using data products derived from the CELEBI pipeline) are required to assess this effect.

### v2

@@ Abstract
With wide-field phased array feed technology, the Australian Square Kilometre Array Pathfinder (ASKAP) is ideally suited to search for seemingly rare radio transient sources that are difficult to discover previous-generation narrow-field telescopes. The Commensal Real-time ASKAP Fast Transient (CRAFT) Survey Science Project has developed instrumentation to continuously search for fast radio transients (duration $\lesssim$ 1 second) with ASKAP, with a particular focus on finding and localising Fast Radio Bursts (FRBs). Since 2018, the CRAFT survey has been searching for FRBs and other fast transients by incoherently adding the intensities received by individual ASKAP antennas, and then correcting for the impact of frequency dispersion on these short-duration signals in the resultant incoherent sum (ICS) in real-time. This low-latency detection enables the triggering of voltage buffers, which facilitates the localisation of the transient source and the study of spectro-polarimetric properties at high time resolution. Here we report the sample of $43$ FRBs discovered in this CRAFT/ICS survey to date. This includes $22$ FRBs that had not previously been reported: $16$ FRBs localised by ASKAP to $\lesssim 1$ arcsec and $6$ FRBs localised to $\sim 10$ arcmin. Of the new arcsecond-localised FRBs, we have identified and characterised host galaxies (and measured redshifts) for $11$. The median of all $30$ measured host redshifts from the survey to date is $z=0.23$. We summarise results from the searches, in particular those contributing to our understanding of the burst progenitors and emission mechanisms, and on the use of bursts as probes of intervening media. We conclude by foreshadowing future FRB surveys with ASKAP using a coherent detection system that is currently being commissioned. This will increase the burst detection rate by a factor of approximately ten and also the distance to which ASKAP can localise FRBs.
@@ 1 INTRODUCTION
Here we summarise the results of this CRAFT/ICS survey and describe the bursts detected. The host galaxies for many of the FRBs discovered are displayed in Figures 1, 2, and 3. In Section 2, we describe the instrumentation and methods used to undertake the searches. We motivate and describe the observing strategies we undertook to find the FRB sample in Section 3 and our multi-wavelength follow up in Section 4. In Section 5, we present the sample of 43 FRBs, of which 37 have measured positions with precisions $\lesssim 1^{\prime\prime}$, of which 30 host galaxies have been identified. We assess survey performance in Section 6, and in Section 7 we provide a review of the main scientific findings and outcomes of the searches. In Section 8, we briefly describe the plan for new FRB search systems for ASKAP.
@@
Figure 1: Montage of FRB host galaxies for ASKAP-localised FRBs with firm associations. We present the FRBs in increasing redshift. In the top left corner, we show the FRB name and redshift. In the top right corner, we list the telescope and observing band of the image. The angular and physical scale at the host redshift are shown in the lower left corner. The 1-$\sigma$ localisation region of the FRB is given by the black ellipse.
@@ 2.1 Hardware
Each PAF receiver comprises $188$ dipole elements arranged in a chequerboard pattern. The signal from each dipole element is transmitted from the PAF to the MRO control building via radio frequency over fibre systems Beresford et al. (2017). Observing is undertaken in one of three bands (bands 1,2, and 3, with 1 being the lowest and 3 the highest). Within each band, it is possible to tune the observing frequencies best suited to the science goals. The signal from each element is then digitised and channelised in a digital receiver Brown et al. (2014). The signals are channelised using a polyphase filter into coarse $1$ MHz channels oversampled by a factor of 32/27 Tuthill et al. (2012).
@@ 2.2 FRB Searches
Despite the MRO having low levels of nearby interference, RFI is present in the searches. The largest sources of interference are satellite interference from global navigation systems in band-2 observations and ionospheric ducting of mobile phone signals (base towers and handsets) in band 1. Data were also affected by interference associated with lightning and electrical storms (even when storms were more than $100$ km from the observatory). Real-time RFI flagging is conducted using a number of heuristics. Frequency channels with abnormally high kurtosis are set to zero, while remaining channels are normalised to a mean of zero and unit standard deviation. Impulsive RFI is removed by subtracting the frequency integrated signal at zero dispersion measure (DM). This is applied dynamically - additionally, a hard-coded list of known RFI channels is maintained, and these frequencies are zeroed for all observations. The methods were developed during commissioning of the ICS system (and the previous fly’s eye survey), with the algorithms tuned to reduce the false positive rate. The fraction of data flagged depended on the observing band, and was typically $10-20\%$. The level of false trigger varied between bands. In the band 1, after flagging there could be cases where there were no false triggers for many days. In band 2, false triggers would occur at a rate of a few per day to one per hour. In the presence of electrical storms, lightning produce false triggers at rate of many per hour. The effectiveness of the mitigation was tested through observations of pulsars. We confirmed that we were detecting the pulses at the expected rate and signal to noise ratio as had been conducted for the fly’s eye survey James et al. (2019b).
@@
The searches produce a stream of raw candidates (time, signal-to-noise ratio, dispersion measure, width) which are broadcast over User Datagram Protocol. A burst (either RFI or a bona-fide FRB) produces candidates over a region of candidate parameter space. A density-based spatial clustering algorithm Ester et al. (1996) is used to group raw candidates. This is important as it allows RFI (which typically has the greatest S/N at low DM) to be distinguished from FRBs. Candidates were clustered using based on time (expressed as sample number within the scan), boxcar width, and dispersion measure (expressed as integer DM trial reported by FREDDA). The density based spatial clustering algorithm used parameters $\epsilon=1$ and $n_{\rm min}=1$ in the clustering. To ensure that the real-time clustering algorithm was working properly, raw candidates were also inspected off-line after being clustered using a friends-of-friends algorithm Huchra & Geller (1982). The friends of friends algorithm clustered based on DM (measured as integer DM units in FREDDA) and time (measured as samples within a scan). The temporal clustering was normally set to a distance of $32$ samples. DM clustering was set to a distance of $20$ integer DM units.
The search latency ranges from a few hundred milliseconds to a few seconds. The time resolution of searches varied over the course of the survey, between $860\,\mu$s and $1.7$ ms, as system performance was optimised. There is a trade-off between the number of antennas that can be ingested and the time resolution. For the majority of the searching, a time resolution of $1.182$ ms was chosen, which allowed the incoherent sum of intensities from $\sim$24 antennas to be formed.
@@ 4.2 Host Associations
We revisit this assumption and suggest an updated formulation for this offset prior. Figure 6 shows the normalised PDF of the angular offsets $\theta$ relative to $\phi$ for 32 FRBs. This FRB sample was restricted to the ICS sample presented in this manuscript with a secure association (see section 5.2) and with $P_{\rm host}>0.90$ using the original PATH analysis and priors. We also include the ASKAP-localised CHIME repeating source FRB 20201124A (Fong et al., 2021). Approximately half of the sample shows $\theta/\phi<1$ with the remainder exhibiting a tail to $\theta/\phi\approx 5$.
Overplotted on the observed distribution are the exponential priors for host galaxy offset with scale lengths of $\phi$ and $\phi/2$. These have been convolved with the reported uncertainties in the FRB localisations. They have also been weighted by a geometrical factor ($2\pi\theta$) which disfavours low offsets [^4] We note that Figure 11 of Aggarwal et al. (2021) failed to include this factor.. Clearly, the exponential with scale-length of $\phi$ is disfavoured while the data are reasonably well-described by the smaller scale length ($\phi/2$). We advocate adopting this new prior for future work on FRBs with PATH, and we utilise it throughout the manuscript. It is also possible as the sample continues to increase that one will adopt a different functional form to better describe the distribution.
Reanalysing all of the FRBs presented in Bhandari et al. (2020b), we find few changes in the posterior probabilities. Most of them were previously $P_{\rm host}>0.95$ and the values increased towards 1. The only notable changes were significant increases in $P_{\rm host}$ for FRB 20181112A and FRB 20191001A. The former FRB is associated with a $z=0.4755$ which has a $z=0.3674$ galaxy in the foreground Prochaska et al. (2019) The latter FRB is associated with a $z=0.2340$ galaxy which is separated by $\approx 5$” from a galaxy at a similar redshift ($z=0.2339$ Bhandari et al., 2020a). In Tables 3 to LABEL:tab:path4 in the Appendix we present PATH probabilities for nearby host galaxies for all ASKAP-localised FRBs.
Figure 6: Comparison of PATH prior distributions (curves) to FRB offsets (histogram) obtained through CRAFT/ICS observations. Here we have restricted to those FRBs with PATH posteriors $P(O|x)$ in excess of 0.9 using the original priors. This corresponds to the gray curve which is an exponential with scale length equal to the half-light radius of the galaxy $\phi$. We now advocate the prior described by the black curve with scale-length equal to $\phi/2$.
@@ 4.3 Optical spectroscopy
We obtained spectroscopic observations of the probable host galaxy candidates that lacked archival data, both to determine the redshift but also to study the global properties of the hosts. These were primarily obtained with FORS2 or X-shooter Vernet et al. (2011) on the VLT, or LRIS or the Deep Imaging Multi-Object Spectrograph (DEIMOS; Faber et al. 2003) on Keck. Brighter host galaxies were observed with the Goodman Spectrograph on the Southern Astrophysical Research Telescope (SOAR, Clemens et al., 2004). Spectroscopic data reduction is described in detail in Glowacki et al. (2023), Gordon et al. (2023), Ryder et al. (2023), and A. Muller et al. (in prep). Uncertainties in the redshifts derived from spectroscopy are approximately $0.0001$ for X-shooter and $0.005$ for other instruments.
@@ 5 FAST RADIO BURST DISCOVERIES
Table 1 in the Appendix summarises the key properties of the FRBs. Table 2 in the Appendix lists the positions of the FRBs, derived interferometrically (where possible) and with the multi-beam localisation method Bannister et al. (2017). Dedispersed dynamic spectra of the bursts from the ICS search data stream, arranged from lowest to highest dispersion measure, can be found in Figures 20 to 24 in the appendix. Even with the relatively low time and frequency resolution of the search data stream the bursts show a variety of spectral and temporal morphologies. Their high-time-resolution properties are reported elsewhere (D. Scott et al., in prep.). We explore the spectral modulation of the bursts in Section 6.9.
@@ 5.2 New localised FRBs
We summarise the properties and localisations of a few other FRBs discovered in the ICS searches, though detailed analyses of the bursts and their host galaxies are deferred to subsequent studies.
@@ 5.2.14 FRB 20240304A
[…]
@@ 6 SURVEY PERFORMANCE
Given the commensal nature of the survey, we assess detection rates as a function of changing observation parameters. Through to the end of 2023 (MJD 60309), ASKAP observed in ICS mode for a total of 15 324.715324.715\,324.715 324.7 hr, for an average on-sky efficiency of $32\%$. As a commensal survey, this has resulted in a variety of pointing directions and frequency configurations. Full logging information was available only for 12 701.712701.712\,701.712 701.7 hr of this, due to an early logging error. This corresponds to the latter 30 FRBs detected until the end of 2023. Furthermore, the current version (‘v3’; 28 FRBs) of the FREDDA FRB detection algorithm has been used only from April 2020 — the initial version (‘v1’; 7 FRBs) had slightly reduced sensitivity to high-DM FRBs, while ‘v2’ (2 FRBs) had a bug that reported incorrect S/N values (Hoffmann et al., 2024). Therefore, some analyses below will be applied to a subset of the data. This will be noted in each case.
@@ 6.2 Modelling frequency dependence
We model the frequency dependence of the detection rate $R(\nu)$ to be
$R(\nu)$ $\propto$ $T(\nu)F(\nu)W(\nu)N_{\rm ant}^{0.75}\nu^{\alpha}\,,$ (1)
@@ 6.4 Source counts analysis
The observed source counts distribution is also a good diagnostic tool for biases in the search pipeline: human vetting, RFI mitigation algorithms, or both, may reject candidates that are close to the nominal detection threshold, or well above it (Macquart & Ekers, 2018b). We have performed Kolmogorov-Smirnov tests (Kolmogorov, 1933; Smirnov, 1948) for consistency with a pure $\alpha=-1.50$ power-law on ICS FRBs, with p-values quoted in Figure 11, and find no strong statistical evidence for a deviation. The same conclusion is reached when dividing the sample into FRBs detected in low, medium, and high frequency ranges.
@@ 6.9 Burst modulation
Many of the ASKAP FRBs show spectral modulation: intensity variations in frequency that could either be diffractive scintillation or intrinsic to the emission mechanism. This was identified in the fly’s eye surveyShannon et al. (2018). We investigated if there is any correlation between modulation and dispersion measure using the FRBs in the search data stream. We choose to investigate modulation and dispersion measure to consider both the fly’s eye and ICS FRBs. We calculate the spectral modulation index to be
@@
where $\sigma_{\rm FRB}^{2}$ is the variance on pulse and $\sigma_{n}^{2}$ is the variance of (thermal) noise measured in a segment of data of equal temporal duration close to the FRB. We estimate the uncertainty on $m_{I}$ using bootstrap Efron (1979). For each FRB, we resample (with replacement) the spectrum and re-calculate $m_{I}$ on each of the realisations. We use the standard deviation of these resample modulation indices as the uncertainty.
While FRBs with DM $\lesssim 500\,$pc cm-3show a variety of spectral modulation, those with DM $>800\,$pc cm-3show little spectral modulation. Spectra of two FRBs with high modulation and two FRBs with high DM are shown in Figure 16. It is possible that spectral modulation is being quenched at higher DM. If scintillation in the Milky Way is causing the spectral modulation, quenching would occur if the bursts are spatially resolved by Milky Way scattering screens, i.e. had been scatter broadened by a second screen. As there is a correlation between dispersion measure and distance Macquart et al. (2020), this would suggest the extra-Galactic scattering scr Studies of a larger sample of bursts in higher spectral and time resolution (such as using data products derived from the CELEBI pipeline) are required to assess this effect. Study of bursts produced by the CELEBI pipeline also benefit from the array-coherennt improvement in signal to noise ratio. This could be complemented with bursts from CHIME/FRB, DSA-110 Sharma et al. (2024), or MeerKAT Jankowski et al. (2023). The latter two would allow for comparison of large FRB samples detected at comparable central frequencies.
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 16: Pulse averaged spectrum of high and low DM FRBs. The spectrum is shown as a black solid line. An off pulse spectrum is shown offset in grey. The dashed horizontal lines show zero fluence. FRB 2020610A (DM $343.8$ pc cm-3) has the highest modulation index of the FRBs in our sample. FRB20180924B (DM $362.4$ pc cm-3) also has high modulation. FRBs 20220610 (DM $1458.1$ pc cm-3) and FRB20210407E (DM $1785.3$ pc cm-3) have the largest DM of the ICS FRB.